ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

4. Algebraic Geometry

1991 MS Classification: 14 Esteves, Eduardo, IMPA, Brazil Degenerations of Weierstrass points In the mid 80’s D. Eisenbud and J. Harris developed the general theory of limit linear series to explain what happens to a linear system and its ramification points on a smooth curve degenerating to a curve of compact type. In one of their papers on this subject they asked: “What are the limits of Weierstrass points in families of curves degenerating to stable curves not of compact type?” I will show how to answer this question first within the general framework of limit linear series, and then directly, assuming generic conditions. Reference: Esteves, Eduardo, Linear systems and ramification points on reducible nodal curves, To appear in Matemática Contemporânea.